bomacharles:

find the probability that a number chosen at random between 200 and 800 inclusive will be a multiple of 9

jhonyy9:

so like a first step what is a number multiple of 9 ? what mean this ? - between 200 and 800 how many the same numbers are there ?

surjithayer:

first find the least number >=200 to do so divide by 9 9| 200


22-2 9-2=7 add to 200 200+7=207 first number is 207 to find last number divide 800 by 9 9|800
88-8 subtract from 800 800-8=792 so numbers are 207,216,223,...,792 find the nummber of numbers.

surjithayer:

it is an A.P. sequence. d=9 a1=207 an=792 use the formula to find n. an=a1+(n-1)d this gives number of favorable events. total number of events=800-200+1=601 (as both 200 and 800 are calculated.) then finally find the probability.